OpenStudy (anonymous):

What is the percent of increase from 25 to 35? A. 10% B. 29% C. 40% D. 71%

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

10 is what percent of 25? 10/25=2/5= 0.4 * 100= 40%

OpenStudy (anonymous):

so the answer is C

OpenStudy (anonymous):

Thank you :)